OpenStudy (anonymous):

solve the system by elimination -5x-y=3 -6x+y=8

OpenStudy (anonymous):

1x + y=11

jimthompson5910 (jim_thompson5910):

Hint: Add the two equations to get -5x-y=3 -6x+y=8 --------- -11x+0y = 11 So -11x = 11

OpenStudy (anonymous):

what about the y

jimthompson5910 (jim_thompson5910):

what did you get for x

OpenStudy (anonymous):

1

jimthompson5910 (jim_thompson5910):

no that's not correct

OpenStudy (anonymous):

can you show me how to get the correct answer

jimthompson5910 (jim_thompson5910):

-11x = 11 how would you isolate x

OpenStudy (anonymous):

I don't know could you show me how to isolate x in numbers though